WayneCorp, Inc., has sales of $555,000, costs of $300,000, depreciation expense of
$40,000, interest expense of $12,000, and a tax rate of 32 percent. (Do not include the dollar sign ($).)
What is this firm's net income?
Answer #1

Calculation of net income:

Earning before interest and tax= sales-cost-depreciation

Earning before interest and tax= 555000-300000-40000= 215000

Earning before tax= earning before interest and tax- interest

Earning before tax= 215000-12000= 203000

Earning after tax= 203000*(1-0.32)= 203000*0.68= 138040

Net income= 138040